# 1]organization=Computer Network Information Center, Chinese Academy of Sciences, addressline=CAS Informatization Plaza No.2 Dong Sheng Nan Lu, Haidian District, city=Beijing, country=China

### v1

@@ 1 Introduction
The strong force, one of the four fundamental forces in nature, is modelled by the theory of quantum chromodynamics (QCD) [1]. Due to the model itself, QCD cannot be solved analytically in some energetic regimes, and in order to obtain predictions under such conditions one has to employ numerical methods via what is known as Lattice QCD (LQCD) [2]. LQCD simulations are amongst the most demanding applications in scientific computing, in terms of High Performance Computing (HPC) usage. Furthermore, one of the most frequent and expensive “fundamental" operations within LQCD simulations is the solution of linear systems of the form $Dx=b$, where $D$ is the Dirac operator. The form of $D$ varies depending on the discretization procedure, leading to different formulations of LQCD, such as Wilson, twisted mass, staggered, and others [2, 3, 4, 5].
@@ 2 Multigrid on ORISE via HIP
In this work we, among other things, translate the current CPU+GPU version of DD-$\alpha$AMG [15] via the Heterogeneous-Compute Interface for Portability (HIP) [16, 17]. The numerical experiments presented in subsect.s 3.1 and 3.2 run on the ORISE supercomputer, hence we start this section by describing the architecture in ORISE. This is followed by a brief introduction to AMG and to the AMG solver DD-$\alpha$AMG, which is being extended here. We then do a comparative description of two iterative methods available in DD-$\alpha$AMG as smoothers: the Generalized Minimal Residual (GMRES) method and Schwarz Alternating Producedure (SAP), discussing in particular their pros and cons in terms of smoothing and of running on CPUs and GPUs; we extend the discussion (and DD-$\alpha$AMG itself) by adding two more iterative methods to the stack of smoothers: Generalized Conjugate Residual (GCR) and modified Richardson. We finalize this section by briefly describing CUDA and HIP, emphasizing their differences and similarities.
@@ 2.1 The ORISE Supercomputer
Direct GPU-to-GPU communication is not possible in ORISE, hence we resort in our implementations to the more expensive procedure of doing GPU-to-CPU followed by CPU-to-GPU in order to be able to communicate data from a GPU to another.
@@ 2.2 Algebraic Multigrid
Whether we are working with geometric [19, 20] or algebraic [21, 22] multigrid depends on what information we use to build $P$ and $R$. In geometric multigrid the structure of the grid is the sole ingredient needed, e.g. one can design $P$ such that it interpolates (i.e. averages) values on lattice sites when going from the coarse to the fine grid, and then one can choose for example $R=P^{H}$. This choice of $R=P^{H}$ is known as Galerkin construction [19].
@@ 2.3 Domain Decomposition Aggregation-Based Algebraic Multigrid
The DD-$\alpha$AMG solver has a setup and a solve phase. During the setup phase it builds the multigrid hierarchy. Let us define this multigrid hierarchy as the set of operators $D_{\ell}$ with $\ell=1,...,L$, and $P_{\ell}$ and $R_{\ell}=P_{\ell}^{H}$ with $\ell=1,...,L-1$, $L$ being the total number of levels. The solve phase then makes use of this hierarchy of operators to solve one or more linear systems i.e. for one or more right hand sides in a sequence.
@@ 3.3 CPU improvements
The time reduction when going from AVX to SSE is larger in tab. 4 than in tab. 3, as expected. This is simply due to having more cache misses in the latter case, where the whole data required by $D_{l}$ cannot be fully stored in L3. It is important to note that AVX-512 usually runs with a reduced frequency compared to AVX2, hence the speedup attained with the former is far from being double than with the latter, matching our results in in tab. 4.
@@
We note from tab.s 3 and 4 the non-uniform speedups due to vectorization as we change the parallelism in the different runs. One way to possibly obtain better and more uniform (i.e. parallelism-independent) vectorization performance is by means of extending the solver to support multiple right hand sides [33]; when applying the Dirac operator on multiple vectors, one has to apply each of the $2N_{v}\times 2N_{v}$ blocks on multiple vectors, which can be made a highly cache-friendly operation.
@@ 5 Acknowledgments
We would like to thank Yibo Yang for providing us with the gauge configuration. Gustavo Ramirez-Hidalgo’s work is supported by the German Research Foundation (DFG) research unit FOR5269 “Future methods for studying confined gluons in QCD”.
@@ Appendix A Configuration and solver parameters
Table 5: Configuration used together with its parameters. Provided by C11P14L[34].

### v2

@@ 1 Introduction
The strong force, one of the four fundamental forces in nature, is modeled by the theory of quantum chromodynamics (QCD) [1]. Due to the model itself, QCD cannot be solved analytically in some energetic regimes, and in order to obtain predictions under such conditions one has to employ numerical methods via what is known as Lattice QCD (LQCD) [2]. LQCD simulations are amongst the most demanding applications in scientific computing, in terms of High Performance Computing (HPC) usage. Furthermore, one of the most frequent and expensive “fundamental" operations within LQCD simulations is the solution of linear systems of the form $Dx=b$, where $D$ is the Dirac operator. The form of $D$ varies depending on the discretization procedure, leading to different formulations of LQCD, such as Wilson, twisted mass, staggered, and others [2, 3, 4, 5].
@@ 2 Multigrid on ORISE via HIP
In this work we, among other things, translate the current CPU+GPU version of DD-$\alpha$AMG [15] via the Heterogeneous-Compute Interface for Portability (HIP) [16, 17]. The numerical experiments presented in subsect.s 3.1 and 3.2 run on the ORISE supercomputer [^1] Although all the final tests have been run on ORISE, part of the code development and testing was done on the supercomputer JUWELS at Jülich Supercomputing Centre., hence we start this section by describing the architecture in ORISE. This is followed by a brief introduction to AMG and to the AMG solver DD-$\alpha$AMG, which is being extended here. We then do a comparative description of two iterative methods available in DD-$\alpha$AMG as smoothers: the Generalized Minimal Residual (GMRES) method and Schwarz Alternating Producedure (SAP), discussing in particular their pros and cons in terms of smoothing and of running on CPUs and GPUs; we extend the discussion (and DD-$\alpha$AMG itself) by adding two more iterative methods to the stack of smoothers: Generalized Conjugate Residual (GCR) and modified Richardson. We finalize this section by briefly describing CUDA and HIP, emphasizing their differences and similarities.
@@ 2.1 The ORISE Supercomputer
Direct GPU-to-GPU communication is not possible in ORISE, hence we resort in our implementations to the more expensive procedure of doing GPU-to-CPU followed by CPU-to-GPU in order to be able to communicate data from one GPU to another.
@@ 2.2 Algebraic Multigrid
Whether we are working with geometric [19, 20] or algebraic [21, 22] multigrid depends on what information we use to build $P$ and $R$. In geometric multigrid, the structure of the grid is the sole ingredient needed, e.g. one can design $P$ such that it interpolates (i.e. averages) values on lattice sites when going from the coarse to the fine grid, and then one can choose for example $R=P^{H}$. This choice of $R=P^{H}$ is known as Galerkin construction [19].
@@ 2.3 Domain Decomposition Aggregation-Based Algebraic Multigrid
The DD-$\alpha$AMG solver has a setup and a solve phase. During the setup phase, it builds the multigrid hierarchy. Let us define this multigrid hierarchy as the set of operators $D_{\ell}$ with $\ell=1,...,L$, and $P_{\ell}$ and $R_{\ell}=P_{\ell}^{H}$ with $\ell=1,...,L-1$, $L$ being the total number of levels. The solve phase then makes use of this hierarchy of operators to solve one or more linear systems i.e. for one or more right-hand sides in a sequence.
@@ 3.3 CPU improvements
The time reduction when going from AVX to SSE is larger in tab. 4 than in tab. 3, as expected. This is simply due to having more cache misses in the latter case, where the whole data required by $D_{l}$ cannot be fully stored in L3. It is important to note that AVX-512 usually runs with a reduced frequency compared to AVX2, hence the speedup attained with the former is far from being double than with the latter, matching our results in tab. 4.
@@
We note from tab.s 3 and 4 the non-uniform speedups due to vectorization as we change the parallelism in the different runs. One way to possibly obtain better and more uniform (i.e. parallelism-independent) vectorization performance is by means of extending the solver to support multiple right-hand sides [33]; when applying the Dirac operator on multiple vectors, one has to apply each of the $2N_{v}\times 2N_{v}$ blocks on multiple vectors, which can be made a highly cache-friendly operation.
@@ 5 Acknowledgments
We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions [34], which are generated on HPC Cluster of ITP-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. Gustavo Ramirez-Hidalgo’s work is supported by the German Research Foundation (DFG) research unit FOR5269 “Future methods for studying confined gluons in QCD”. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC), under the project with id MUL-TRA.
@@ Appendix A Configuration and solver parameters
Table 5: Configuration used together with its parameters provided by the CLQCD collaboration. For more details of the ensembles, we refer to Refs.[34][35].